OpenStudy (anonymous):

Differentiate 3yx^2=x^3

OpenStudy (anonymous):

implicit diff right? wrt x \[3(2xy+x^2y')=3x^2\]solve for y' algebra from here on in

OpenStudy (anonymous):

\[2xy+x^2y'=x^2\] \[x^2y'=x^2-2xy\] \[y'=\frac{x^2-2xy}{x^2}=\frac{x-2y}{x}\] check my algebra because it is way early

OpenStudy (anonymous):

Good work.